Parth (parthkohli):

Given: \(a + b = 4\) \(a^3 + b^3 = 28\) What is \(a - b\)? I got this: \(a^3 + b^3 = (a + b)(a^2 - ab + b^2)\) \(28 = 4(a^2 - ab + b^2)\) How do I continue?

OpenStudy (lgbasallote):

you can simplify it further

Parth (parthkohli):

\(7 = a^2 - ab + b^2\)

Parth (parthkohli):

I completed it by guess and check, but I need to do it algebraically

OpenStudy (lgbasallote):

\[a^2 - ab + b^2 = (a+b)^2 - 3ab\]

Parth (parthkohli):

@Waterineyes, (a - b)^2*

OpenStudy (lgbasallote):

@waterineyes \[a^2 + b^2 -ab - ab = (a-b)^2 \ne (a+b)^2\]

mathslover (mathslover):

\[\large{(a-b)^2=a^2+b^2-2ab}\] \[\large{a^2+b^2+2ab=16}\] \[\large{(a-b)^2=a^2+b^2+2ab-4ab=16-4ab}\]

OpenStudy (anonymous):

\[(a + b)^2 - ab - ab =a^2 + b^2\]

mathslover (mathslover):